Corporate Finance

nDuring the year, Bigquzi Co. had sales of 3,000,000; cost of goods sold, S&A and depreciation were 1,600,000,450000,540000 respectively. In addition the company had an interest expense of 250,000 and a tax rate of 35%. (Ignore any tax loss carryback or carryforward provisions)


1)What is Bigquzi’ net income and operating cash flow

2) The company paid out 400000, is this possible without issuing additional equity/debt? 

3) If net capital spending and investment in net working capital is zero, and the firm issue no new equity, what do you know about the firm’s long-term debt account?

4) If the company decides to spend $200000 on a new project, what happens to the firm’s debt account?
Answer #1

1. Net income = sales cost of goods sold-sales, general and administrative expenses-depreciation-interest expenses** (1-tax rate) = (3,000,000-1,600,000-450,000-540,000-250,000) * (1-35%) = $104,000


Operating cash flow = net income + depreciation = 104,000 + 540,000 = $644,000


2. No need to issue any debt/equity, because the net income is $644,000, which is enough to pay a dividend of $400,000.


3. If there is no capital and working capital investment, it will not have much impact on the long-term debt account. The long-term debt account decreases in value due to the repayment of the principal during the period.


4. The debt account increases in proportion to the capital structure.


answered by: gavin